OpenStudy (anonymous):

my teacher asked me what's 1+1-1+1-1+1-1, What's the answer?

OpenStudy (anonymous):

http://lmgtfy.com/?q=1%2B1-1%2B1-1%2B1-1

OpenStudy (anonymous):

its one just do the problem in steps one at a time, and you will get 1. 1+1=2-1=1+1=2-1=1+1=2-1=1

OpenStudy (anonymous):

1

OpenStudy (beccaboo333):

Welcome to OpenStudy, first off this is in the wrong section. You should also read the CoC. https://openstudy.com/code-of-conduct